1 Quadrature Rule A quadrature rule is a way to approximate integrals numerically i.e. using a computer). Many such quadraturExplain using Matlab code but also why you used the linear system please
Answer #1

5124t 1- o2/6 3 + =ON+ O oo16 t o 2 56 S. + 0.0777 6 w+ o.327 68 wp +/04 per To solve thi-s systen Matlab

% matlab programe

x=0:0.2:1;% vector of x values

A=[[1,1,1,1,1,1];x;x.^2;x.^3;x.^4;x.^5]% coefficient matrix of the linear %system
B=[1;1/2;1/3;1/4;1/5;1/6]; % right hand side vector of the system
w=inverse(A)*B; % where w = [w0, w1, w2, w3, w4, w5]

% output

w =

   0.065972
   0.260417
   0.173611
   0.173611
   0.260417
   0.065972

% w0 =0.065972, w1= 0.260417 , w2 = 0.173611, w3 = 0.173611 , w4 =0.260417 , w5 =0.065972